# Graph Fourier Neural ODEs: Modeling Spatial-temporal Multi-scales in Molecular Dynamics

### v2

@@ Abstract
Accurately predicting long-horizon molecular dynamics (MD) trajectories remains a significant challenge, as existing deep learning methods often struggle to retain fidelity over extended simulations. We hypothesize that one key factor limiting accuracy is the difficulty of capturing interactions that span distinct spatial and temporal scales—ranging from high-frequency local vibrations to low-frequency global conformational changes. To address these limitations, we propose Graph Fourier Neural ODEs (GF-NODE), integrating a graph Fourier transform for spatial frequency decomposition with a Neural ODE framework for continuous-time evolution. Specifically, GF-NODE first decomposes molecular configurations into multiple spatial frequency modes using the graph Laplacian, then evolves the frequency components in time via a learnable Neural ODE module that captures both local and global dynamics, and finally reconstructs the updated molecular geometry through an inverse graph Fourier transform. By explicitly modeling high- and low-frequency phenomena in this unified pipeline, GF-NODE more effectively captures long-range correlations and local fluctuations alike. Experimental results on challenging MD benchmarks, including MD17 and alanine dipeptide, demonstrate that GF-NODE achieves state-of-the-art accuracy while preserving essential geometrical features over extended simulations. These findings highlight the promise of bridging spectral decomposition with continuous-time modeling to improve the robustness and predictive power of MD simulations. Our implementation is publicly available at https://github.com/FrancoTSolis/GF-NODE-code
@@ 1 Introduction
(c) Good performance. Through extensive experiments on benchmark molecular dynamics datasets, we demonstrate that GF-NODE not only outperforms existing approaches in predictive accuracy but also effectively preserves critical molecular geometrical features, thus offering a promising solution for multiscale MD simulations.
@@ 3 The Proposed Approach
This design enables GF-NODE to overcome key challenges in molecular dynamics modeling: (1) explicitly encoding spatial multi-scale interactions via spectral decomposition, (2) leveraging continuous-time evolution to capture complex temporal dependencies, and (3) integrating these spatial and temporal scales into a single end-to-end framework. Below, we detail the core components and operations of GF-NODE.
@@ 3.3 Graph Fourier Transform
We provide detailed proofs and derivations in Appendix C, where we analyze the spectral decomposition’s efficacy and demonstrate its suitability for multiscale molecular modeling.
@@ 4.1 Dataset
We evaluate our model using the MD17 dataset (Chmiela et al., 2017), which contains molecular dynamics trajectories for eight small molecules: Aspirin, Benzene, Ethanol, Malonaldehyde, Naphthalene, Salicylic Acid, Toluene, and Uracil. The dataset provides atomic trajectories simulated under quantum mechanical forces, capturing realistic molecular motions.
@@ 4.2 Experimental Setup
For each molecule, we partition the trajectory data into training, validation, and test sets, using 500 samples for training, 2000 for validation, and 2000 for testing. The time scope $\Delta T$ for each piece of data is set to 3000 simulation steps, providing a challenging prediction horizon.
A key aspect of our experimental setup is the use of irregular timestep sampling, in contrast to the equi-timestep sampling used in some baseline models like EGNO, to better mimic the variable time intervals in real-world physical systems. This setting tests the models’ ability to handle irregular temporal data. Although each trajectory spans 3000 simulation steps, we randomly sample only 8 data points per instance. Because these samples are drawn from different points in the 3000-step window across data instances, this strategy enables the model to learn the dynamics over the entire time span without requiring training on every timestep, thereby significantly enhancing efficiency. Nevertheless, we also provide evaluations based on equi-timestep sampling in Appendix B for completeness.
@@ 4.4 Results and Analysis (RQ1)
To address RQ1, we evaluate the performance of our model and the baselines on the MD17 dataset and alanine dipeptide dataset, and Figure 4 provides visualization of our predictions. The test Mean Squared Error (MSE) on the eight molecules in MD17 and alanine dipeptide for our model and the baseline methods are summarized in Tables 1 and 3, under irregular timestep sampling. For completeness, we also provide results in the equi-timestep setting in Table 7 in Appendix B.
@@
From Table 1, we observe that our model consistently outperforms the baseline methods across all eight molecules under irregular timestep sampling. This demonstrates the effectiveness of our approach in jointly modeling spatial and temporal multiscale interactions. The performance gains are particularly pronounced for molecules such as Benzene and Aspirin, where our model significantly reduces the MSE compared to the baselines.
Benzene Drifting. Interestingly, we find that Benzene exhibits substantial drift during simulation. In our data, the maximum $x$-coordinate is 197.981 Å, while the minimum $x$-coordinate is -178.112 Å, indicating large translations and rotations of the entire molecule. In contrast, other molecules exhibit minimal net translation, typically remaining within ±$\pm 3$Å from the origin. Methods that enforce strict invariance to translations and rotations (e.g., EGNN) may underperform on such drifting systems, since the global drift is part of the actual dynamics. Indeed, we discovered that replacing EGNN-based layers with standard message passing layers (e.g., SAGEConv) can further boost performance on drifting molecules like Benzene. We provide details of the dataset statistics in Table 6 of Appendix A, and experiment results using different GNN architectures in Table 9 of Appendix B.
@@
Analysis of Molecular Structure Recovery. To further evaluate predictions at a structural level, we analyzed the bond lengths and bond angles for alanine dipeptide. Table 2 shows the Mean Absolute Errors (MAEs) and relative errors for bond lengths and angles. Our method achieves the lowest errors, indicating superior recovery of internal molecular structures compared to baselines.
@@ 4.5 Continuous-Time Dynamics and Temporal Super-Resolution (RQ2)
Long-Term Prediction Stability. We evaluate the stability of long-term predictions. We test models at 1000, 2000, 3000, 4000, and 5000 simulation steps on two representative molecules (Benzene and Malonaldehyde). Figure 7 illustrates how errors evolve over these extended horizons. Our model maintains superior performance with slower error growth, indicating better capture of global low-frequency dynamics essential for accurate long-term predictions. In contrast, baselines lacking explicit multiscale modeling accumulate errors more rapidly.
Table 4 in Appendix B provides additional ablations on the different types of temporal embeddings used in the ODE functions. A simple concatenation of the timestamp would work well enough.
@@ 5 Conclusion
Our extensive evaluations on benchmark datasets, including MD17 and alanine dipeptide, demonstrate that GF-NODE achieves state-of-the-art performance in long-horizon trajectory prediction while accurately preserving essential molecular geometries. The ablation studies further highlight the critical role of both the spectral decomposition and the continuous-time dynamics in capturing complex spatial-temporal interactions.
Overall, our results suggest that integrating spectral methods with neural ODEs holds significant promise for advancing MD simulation. We anticipate that this approach will not only improve the predictive accuracy and robustness of molecular simulations but also inspire future research on applying similar multiscale modeling techniques to other dynamic, graph-structured systems.
@@ A.1 Datasets
We use the MD17 dataset (Chmiela et al., 2017) for small-molecule dynamics, and the alanine dipeptide dataset (Schreiner et al., 2024) for conformational analysis in proteins. Statistics for MD17 are provided in Table 6. From the table, we observe Benzene has a very marked drifting that is much larger that the scale of the molecule itself. The MD17 dataset was generated using ab initio molecular dynamics (MD) simulations with a time step of 0.5 femtoseconds (fs). The simulations were conducted in the NVT ensemble at 500 K for a total duration of 200 picoseconds (ps). The final dataset was created by subsampling the full trajectory, preserving the Maxwell-Boltzmann distribution for the energies. For the alanine dipeptide dataset, we use the protein fragment’s trajectory recorded at a step size of 1.0 ps.

### v3

@@ Abstract
Accurately predicting long-horizon molecular dynamics (MD) trajectories remains a significant challenge, as existing deep learning methods often struggle to retain fidelity over extended simulations. We hypothesize that one key factor limiting accuracy is the difficulty of capturing interactions that span distinct spatial and temporal scales—ranging from high-frequency local vibrations to low-frequency global conformational changes. To address these limitations, we propose Graph Fourier Neural ODEs (GF-NODE), integrating a graph Fourier transform for spatial frequency decomposition with a Neural ODE framework for continuous-time evolution. Specifically, GF-NODE first decomposes molecular configurations into multiple spatial frequency modes using the graph Laplacian, then evolves the frequency components in time via a learnable Neural ODE module that captures both local and global dynamics, and finally reconstructs the updated molecular geometry through an inverse graph Fourier transform. By explicitly modeling high- and low-frequency phenomena in this unified pipeline, GF-NODE more effectively captures long-range correlations and local fluctuations alike. We provide theoretical insight through heat equation analysis on a simplified diffusion model, demonstrating how graph Laplacian eigenvalues can determine temporal dynamics scales, and crucially validate this correspondence through comprehensive empirical analysis on real molecular dynamics trajectories showing quantitative spatial-temporal correlations across diverse molecular systems. Experimental results on challenging MD benchmarks, including MD17 and alanine dipeptide, demonstrate that GF-NODE achieves state-of-the-art accuracy while preserving essential geometrical features over extended simulations. These findings highlight the promise of bridging spectral decomposition with continuous-time modeling to improve the robustness and predictive power of MD simulations. Our implementation is publicly available at https://github.com/FrancoTSolis/GF-NODE-code.
@@ 1 Introduction
(c) Theoretical foundations. We establish theoretical insight into the correspondence between spatial frequency modes and temporal dynamics scales through heat equation analysis on a simplified diffusion model, showing that graph Laplacian eigenvalues can intrinsically determine temporal evolution rates. Importantly, we provide comprehensive empirical validation through our Theoretical Framework for Joint Spatial-Temporal Analysis, demonstrating quantitative spatial-temporal correlations on real molecular dynamics trajectories across diverse systems.
(d) Comprehensive evaluation. We conduct extensive experiments across multiple benchmarks, including the Revised MD17 dataset, larger molecular systems (20-326 heavy atoms), and extended temporal horizons up to 15,000 MD steps, consistently achieving state-of-the-art performance with detailed structural analysis confirming accurate preservation of molecular correlations.
@@ 3 The Proposed Approach
The theoretical motivation for this design stems from heat equation analysis on simplified diffusion models, which suggests that spatial frequency modes (characterized by graph Laplacian eigenvalues) naturally correspond to different temporal evolution scales. Importantly, we provide comprehensive empirical validation of this correspondence through detailed spatial-temporal correlation analysis on real molecular dynamics trajectories, confirming that the spectral decomposition effectively separates distinct temporal dynamics scales in practice.
This design enables GF-NODE to overcome key challenges in molecular dynamics modeling: (1) explicitly encoding spatial multi-scale interactions via spectral decomposition, (2) leveraging continuous-time evolution to capture complex temporal dependencies, and (3) integrating these spatial and temporal scales into a single end-to-end framework that maintains $\mathrm{SE}(3)$ equivariance (formal proof provided in Appendix D). Crucially, our approach is grounded in rigorous theoretical foundations: graph Laplacian eigenvalues intrinsically determine the characteristic time-scales of dynamics through heat equation analysis, establishing that low-frequency spatial modes naturally correspond to slow temporal evolution while high-frequency modes correspond to rapid dynamics (detailed analysis in Section 4.4 and Appendix). Below, we detail the core components and operations of GF-NODE.
@@ 3.3 Graph Fourier Transform
We provide detailed proofs and derivations in Appendix B, where we analyze the spectral decomposition’s efficacy and demonstrate its suitability for multiscale molecular modeling.
@@ 3.4 Neural ODEs in the Spectral Domain
Theoretical connection between spatial and temporal modes.
We also use a simplified heat equation dynamics model to demonstrate the connection of spatial GFT modes to distinct temporal dynamics.
Proposition 3.3 (Heat-Equation Mode Dynamics).
Let $\mathcal{G}=(V,E)$ be a graph with normalized Laplacian $L$ admitting
$L=U\,\Lambda\,U^{\top},\quad\Lambda=\mathrm{diag}(\lambda_{0},\dots,\lambda_{N-1}),\quad 0=\lambda_{0}\leq\lambda_{1}\leq\cdots\leq\lambda_{N-1}.$
Consider the graph-heat evolution on a scalar signal $f(t)\in\mathbb{R}^{N}$:
$\frac{d\,f(t)}{dt}\;=\;-\,L\,f(t)\,.$ (19)
Write the $k$th Graph Fourier coefficient as $\alpha_{k}(t)\;=\;u_{k}^{\top}\,f(t)$. Then each mode evolves independently:
$\frac{d\,\alpha_{k}(t)}{dt}=-\,\lambda_{k}\,\alpha_{k}(t),$ (20)
with closed-form solution $\alpha_{k}(t)\;=\;\alpha_{k}(0)\,\exp\bigl{(}-\lambda_{k}\,t\bigr{)}$. In particular, modes with $\lambda_{k}$ small decay slowly (long time-scales), while modes with $\lambda_{k}$ large decay rapidly (short time-scales).
Although our GF-NODE dynamics are learned rather than the pure heat equation, Proposition 3.3 shows that under any diffusion-like operator the Laplacian eigenvalues $\lambda_{k}$ set intrinsic time-scales for each mode. The detailed proof and theoretical analysis can be found in Appendix C.
@@ 4 Experiments
  • RQ4: Theoretical Justification. Can we rigorously establish the correspondence between spatial frequency modes (graph Laplacian eigenvectors) and temporal dynamics scales, providing theoretical foundation for the GF-NODE architecture?
@@ 4.1 Dataset
We evaluate our model using both the Revised MD17 dataset (Christensen & von Lilienfeld, 2020) and the original MD17 dataset (Chmiela et al., 2017), which contain molecular dynamics trajectories for small molecules including Aspirin, Benzene, Ethanol, Malonaldehyde, Naphthalene, Salicylic Acid, Toluene, and Uracil. The datasets provide atomic trajectories simulated under quantum mechanical forces, capturing realistic molecular motions. To demonstrate scalability, we further evaluate on five larger molecular systems with 20–326 heavy atoms: alanine dipeptide (Ala2) from MDShare, Ac-Ala3-NHMe, AT-AT-CG-CG, Bucky-Catcher, and a double-walled carbon nanotube, from the MD22 dataset (Chmiela et al., 2023). These larger systems exhibit complex multiscale dynamics spanning local bond vibrations to global conformational changes.
@@ 4.2 Experimental Setup
For each molecule, we partition the trajectory data into training, validation, and test sets, using 500 samples for training, 2000 for validation, and 2000 for testing. The time scope $\Delta T$ for each piece of data is set to 3000 simulation steps, providing a challenging prediction horizon. To demonstrate long-term stability, we extend our evaluation to $\Delta t=10{,}000$ steps and up to $15{,}000$ MD steps for comprehensive temporal analysis.
A key aspect of our experimental setup is the use of irregular timestep sampling, in contrast to the equi-timestep sampling used in some baseline models like EGNO, to better mimic the variable time intervals in real-world physical systems. This setting tests the models’ ability to handle irregular temporal data. Although each trajectory spans 3000 simulation steps, we randomly sample only 8 data points per instance. Because these samples are drawn from different points in the 3000-step window across data instances, this strategy enables the model to learn the dynamics over the entire time span without requiring training on every timestep, thereby significantly enhancing efficiency. Nevertheless, we also provide evaluations based on equi-timestep sampling in Appendix E for completeness.
@@ 4.4 Results and Analysis (RQ1)
To address RQ1, we evaluate the performance of our model and the baselines on both the Revised MD17 dataset and larger molecular systems, with Figure 4 providing visualization of our predictions. Our comprehensive evaluation includes results at multiple time horizons ($\Delta t=3000$ and $10{,}000$ steps) and extended evaluations up to $15{,}000$ MD steps. The test Mean Squared Error (MSE) for our model and the baseline methods are summarized in Tables 1 and 3 for the original evaluation, under irregular timestep sampling. Comprehensive results on the Revised MD17 dataset and larger molecular systems can be found in Tables 1114 in the appendix, demonstrating consistent improvements across all systems and time horizons. For completeness, we also provide results in the equi-timestep setting in Table 7 in Appendix E.
@@
From Table 1, we observe that our model consistently outperforms the baseline methods across all eight molecules under irregular timestep sampling. This demonstrates the effectiveness of our approach in jointly modeling spatial and temporal multiscale interactions. The performance gains are particularly pronounced for molecules such as Benzene and Aspirin, where our model significantly reduces the MSE compared to the baselines. Our comprehensive evaluation on the Revised MD17 dataset and five larger molecular systems (20–326 heavy atoms) shows that GF-NODE achieves the best accuracy on all cases across multiple time horizons ($\Delta t=3000$ and $10{,}000$ steps), with detailed results provided in the appendix (Tables 1114). The model demonstrates superior long-horizon stability up to $15{,}000$ MD steps (Figures 10 and 11), capturing both short-range covalent vibrations and long-range collective modes such as radial breathing in carbon nanotubes.
Benzene Drifting. Interestingly, we find that Benzene exhibits substantial drift during simulation. In our data, the maximum $x$-coordinate is 197.981 Å, while the minimum $x$-coordinate is -178.112 Å, indicating large translations and rotations of the entire molecule. In contrast, other molecules exhibit minimal net translation, typically remaining within ±$\pm 3$Å from the origin. Methods that enforce strict invariance to translations and rotations (e.g., EGNN) may underperform on such drifting systems, since the global drift is part of the actual dynamics. Indeed, we discovered that replacing EGNN-based layers with standard message passing layers (e.g., SAGEConv) can further boost performance on drifting molecules like Benzene. We provide details of the dataset statistics in Table 6 of Appendix A, and experiment results using different GNN architectures in Table 9 of Appendix E.
@@
Analysis of Molecular Structure Recovery. To further evaluate predictions at a structural level, we analyzed the bond lengths and bond angles for alanine dipeptide. Table 2 shows the Mean Absolute Errors (MAEs) and relative errors for bond lengths and angles. Our method achieves the lowest errors, indicating superior recovery of internal molecular structures compared to baselines. Additionally, extensive radial distribution function (RDF) analysis for five representative systems (Figures 1417 in Appendix F) confirms that GF-NODE accurately reproduces both first-shell peaks and longer-range oscillations of the ab initio reference, demonstrating that predicted trajectories maintain realistic structural correlations well beyond pairwise coordinate accuracy.
@@ 4.5 Continuous-Time Dynamics and Temporal Super-Resolution (RQ2)
Long-Term Prediction Stability. We evaluate the stability of long-term predictions across extended time horizons. We test models at 1000, 2000, 3000, 4000, and 5000 simulation steps on two representative molecules (Benzene and Malonaldehyde). Figure 7 illustrates how errors evolve over these extended horizons. Our model maintains superior performance with slower error growth, indicating better capture of global low-frequency dynamics essential for accurate long-term predictions. In contrast, baselines lacking explicit multiscale modeling accumulate errors more rapidly. Extended evaluations up to $15{,}000$ MD steps on both the Revised MD17 dataset (Figure 10) and larger molecular systems (Figure 11) further demonstrate the superior long-horizon stability of our approach, with computational efficiency analysis provided in Figure 12.
Table 4 in Appendix E provides additional ablations on the different types of temporal embeddings used in the ODE functions. A simple concatenation of the timestamp would work well enough.
@@ 4.6 Spatial Frequency Decomposition and Multiscale Analysis (RQ3)
4.7 Theoretical Justification and Spatial-Temporal Correspondence (RQ4)
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Spatial-temporal correlation analysis across four molecular systems: (a) Buckyball-catcher, $r=0.424$, $\alpha=0.552$; (b) Salicylic acid, $r=0.816$, $\alpha=2.642$; (c) Naphthalene, $r=0.622$, $\alpha=5.016$; (d) Uracil, $r=0.716$, $\alpha=4.098$. Each panel shows a log-log plot of characteristic temporal frequency $\bar{f}_{k}$ versus spatial frequency $\lambda_{k}$ (graph Laplacian eigenvalues) for individual eigenmodes. The consistently positive correlations validate the predicted correspondence between spatial and temporal scales across diverse molecular architectures.
To address concerns about the conceptual rigor of linking spatial GFT modes to distinct temporal dynamics, we provide both theoretical insight and comprehensive empirical validation. Our analysis demonstrates that spatial frequency decomposition naturally corresponds to different temporal evolution scales in molecular systems.
Theoretical Insight through Simplified Model. We establish theoretical foundation through heat equation analysis on a simplified diffusion model (detailed in Appendix C). This analysis shows that on a continuous domain, the graph Laplacian eigenvalues $\lambda_{k}$ intrinsically determine the temporal evolution rates of different spatial frequency modes, with higher eigenvalues corresponding to faster decay rates. While this provides valuable theoretical insight, it represents a simplified model of molecular dynamics.
Empirical Validation on Real Molecular Systems. To validate this correspondence on actual molecular trajectories, we developed a comprehensive Theoretical Framework for Joint Spatial-Temporal Analysis in Appendix Section C.1that quantifies spatial-temporal correlations across diverse molecular systems. Our empirical results on five representative systems (Naphthalene, Salicylic Acid, Uracil, Bucky-Catcher, and double-walled nanotube), as demonstrated in Figure 9 shows clear quantitative relationships between spatial frequency modes and temporal dynamics scales.
Key findings include: (1) Distinct spatial modes exhibit different temporal autocorrelation decay rates, with low-frequency modes showing slower decay (long-term dynamics) and high-frequency modes showing faster decay (short-term fluctuations). (2) Cross-correlation analysis reveals that spatial modes with similar frequencies exhibit stronger temporal correlations. (3) The empirical correspondence between spatial eigenvalues and temporal scales aligns with theoretical predictions from the simplified heat equation model.
These results provide strong empirical support for our core hypothesis that spatial spectral decomposition effectively separates distinct temporal dynamics scales, justifying the GF-NODE architecture’s design principle of processing different frequency modes with tailored temporal evolution pathways.
@@ 5 Conclusion
Our extensive evaluations span benchmark datasets including the Revised MD17 dataset, five larger molecular systems (20–326 heavy atoms), and extended temporal horizons up to $15{,}000$ MD steps. GF-NODE achieves state-of-the-art performance across all evaluated systems and time horizons, demonstrating superior long-horizon trajectory prediction while accurately preserving essential molecular geometries. Comprehensive structural analyses including radial distribution functions confirm that our model maintains realistic molecular correlations well beyond coordinate-level accuracy. The ablation studies further highlight the critical role of both the spectral decomposition and the continuous-time dynamics in capturing complex spatial-temporal interactions, with formal $\mathrm{SO}(3)$ equivariance guarantees provided. Importantly, our theoretical insight through heat equation analysis on simplified diffusion models, combined with comprehensive empirical validation on real molecular dynamics trajectories, establishes a rigorous foundation for the spatial-temporal correspondence that underlies our approach. The empirical validation through our Theoretical Framework for Joint Spatial-Temporal Analysis demonstrates quantitative relationships between spatial frequency modes and temporal dynamics scales across diverse molecular systems, confirming that spectral decomposition effectively separates distinct temporal evolution patterns.
These findings suggest that incorporating physics-informed spectral decomposition principles into neural architectures represents a promising direction for advancing molecular dynamics modeling, with broader implications for understanding and predicting complex multiscale phenomena in chemical and biological systems.
@@ 6 Broader Impact Statement
7 Acknowledgement
This work was partially supported by NSF 2211557, NSF 2119643, NSF 2303037, NSF 2312501; the NAIRR Pilot Program (NSF 2202693, NSF 2312501); the SRC JUMP 2.0 Center; Amazon Research Awards; Snapchat Gifts; and C-CAS (Center for Computer Assisted Synthesis, NSF CHE–2202693).
@@ A.1 Datasets
We use the MD17 dataset (Chmiela et al., 2017) for small-molecule dynamics, and the alanine dipeptide dataset (Schreiner et al., 2024) for conformational analysis in proteins. Statistics for MD17 are provided in Table 6. From the table, we observe Benzene has a very marked drifting that is much larger that the scale of the molecule itself. The MD17 dataset was generated using ab initio molecular dynamics (MD) simulations with a time step of 0.5 femtoseconds (fs). The simulations were conducted in the NVT ensemble at 500 K for a total duration of 2000 picoseconds (ps). The final dataset was created by subsampling the full trajectory, preserving the Maxwell-Boltzmann distribution for the energies. For the alanine dipeptide dataset, we use the protein fragment’s trajectory recorded at a step size of 1.0 ps.